OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2. y = 2x - 8 y = 2x - 4 y = 2x + 7 y = 2x + 5

OpenStudy (texaschic101):

we will be using y = mx + b and solving for b slope(m) = 2 (3,-2) x = 3 and y = -2 now we sub y = mx + b -2 = 2(3) + b -2 = 6 + b -2 - 6 = b -8 = b therefore, your equation is : y = 2x - 8